# Critical point representation of the mutual information in the sparse stochastic block model

### v1

@@ 6 The disassortative and small-time settings
If $C<+\infty$ is as in the statement of Proposition 1.3, and $f:[0,C^{-1}]\times\mathcal{M}_{+}\to\mathbb{R}$ denotes the limit of the sequence of enriched free energies $(\overline{F}_{N})_{N\geq 1}$ on $[0,C^{-1}]\times\mathcal{M}_{+}$, then the map $t\mapsto f(t,0)$ is differentiable at $t=0$.
@@ 7 Critical point selection
where $x\in\Sigma_{N}$ encodes the allocation of students, $y\in\Sigma_{N}$ denotes the allocation of tutors, the label $+1$ represents the group preferring in-person instruction, and the label $-1$ represents the group preferring virtual instruction. We will assume that the labels $X_{i}\sim P^{X}$ and $Y_{i}\sim P^{Y}$ are taken to be i.i.d. from two probability measures $P^{X}$ and $P^{Y}$ on $\Sigma_{1}$. This means that the vector $(X,Y)$ is sampled from the product distribution
@@
[…]
@@
The strategy will be to show that the time derivative of the limit free energy at the origin, which can be shown to exist through similar arguments to those in the proof of Corollary 6.3, does not agree with the time derivative implied by the variational formula (7.23). Determining the time derivative of the enriched free energy (7.17) can be done through a similar computation to that in [18, Lemma 2.1], and reveals that for all $(t,\mu_{1},\mu_{2})\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}^{2}$,
@@
$\partial_{t}f(0,0,0)\geq D_{\mu}\psi_{Y}(0;\delta_{\overline{m}})=\operatorname{\mathbb{E}}\langle c+\Delta\epsilon\overline{m}\rangle_{*}\log\langle c+\Delta\epsilon\overline{m}\rangle_{*}-c-\Delta\overline{m}^{2}.$

### v2

@@ 6 The disassortative and small-time settings
If $C<+\infty$ is as in the statement of Proposition 1.3, and $f:[0,C^{-1}]\times\mathcal{M}_{+}\to\mathbb{R}$ denotes the limit of the sequence of enriched free energies $(\overline{F}_{N})_{N\geq 1}$ on $0,C^{-1}]\times\mathcal{M}_{+}$, then the map $t\mapsto f(t,0)$ is differentiable at $t=0$.
@@ 7 Critical point selection
where $x\in\Sigma_{N}$ encodes the allocation of students, $y\in\Sigma_{N}$ denotes the allocation of tutors, the label +1+1 represents the group preferring in-person instruction, and the label 1-1 represents the group preferring virtual instruction. We will assume that the labels $X_{i}\sim P^{X}$ and $Y_{i}\sim P^{Y}$ are taken to be i.i.d. from two probability measures $P^{X}$ and $P^{Y}$ on $\Sigma_{1}$. This means that the vector $(X,Y)$ is sampled from the product distribution
@@
where the randomness in these two datasets is taken to be independent conditionally on $(X,Y)$. As usual, we write $\langle\cdot\rangle$ for the Gibbs average associated with the Hamiltonian (7.16). This means that for any bounded and measurable function $f=f(x^{1},y^{1}\ldots,x^{n},y^{n})$ of finitely many replicas,
@@
The strategy will be to show that the time derivative of the limit free energy at the origin, which can be shown to exist through similar arguments to those in the proof of Corollary 6.3, does not agree with the time derivative implied by the variational formula (7.23). Determining the time derivative of the enriched free energy (7.17) can be done through a similar computation to that in [19, Lemma 2.1], and reveals that for all $(t,\mu_{1},\mu_{2})\in\mathbb{R}_{\geq 0}\times\mathcal{M}_{+}^{2}$,
@@
$\partial_{t}f(0,0,0)\geq D_{\mu}\psi_{Y}(0;\delta_{\overline{m}})=\operatorname{\mathbb{E}}\langle c+\Delta\varepsilon\overline{m}\rangle_{*}\log\langle c+\Delta\varepsilon\overline{m}\rangle_{*}-c-\Delta\overline{m}^{2}.$